OpenStudy (anonymous):

Check whether (– 1, 5) is a solution of each system of equations. x+y=4 x-y=6

OpenStudy (anonymous):

-1 is our 'x' coordinate, and 5 is our y coordinate. Another way to state this is: (x = -1), (y = 5). All ya have to do now is plug these values into each given equation and then determine whether or not the answer is true. Here. Lemme show ya: x + y = 4 (-1) + (5) = 4 4 = 4. TRUE! So we know that (-1, 5) is a solution of the equation: x + y = 4. Let's try this same procedure with the second equation. x - y = 6 (-1) - (5) = 6 But -1 - 5 comes up to -6. So we end up saying: -6 = 6. Is that true at all! Nopety-nopety-no. So the first equation checks out to be correct, and the second one doesn't! Good luck! Rock on!